4. Categorical data: (from Samuels, Witmer, Schaffner, Statistics for the Life Sciences, 5th edition, page 53). In an effort to determine if there are differenes in the primary sources of fecal contamination at different locations in the Morro Bay watershed, n-623 water specimens were collected at three primary locations that feed into Morro Bay:... DNA fingerprinting techniques were used to determine the intestinal origin of the dominant E coli strain in each water specimen. A frequency table is given below E. coli source Location Bird Domestic Farm Human 38 63 60 Terrestrial Total mammal animal 106 et Chorro Creek Los Osos 79 Creek Baywood35 Seeps Total 241 256 126 623 36 29 56 23 108 32 26 160 138 56 a. What percent of the sources are bird, domestic pet, farm animal, human, and terrestrial mammal? b. c. Which location has the highest rate of contamination from farm animals? Compute d. Which location has the highest rate of contamination from humans? Compute the e. For Los Osos Creek, report the rates of contamination by source in decreasing order f. What percent of the samples with terrestrial mammal E. coli contamination are in What percent of the samples were taken in the three locations? the three rates three rates of rate Chorro Creek? Compute the three rates.
Answer #1

a.

The percentages of sources to be calculated by finding the ratio between the respective totals and the species total.

Percent of the sources are bird = 160 / 623 = 25.68%

Percent of the sources are domestic pet = 108 / 623 = 17.33%

Percent of the sources are farm animal = 138 / 623 = 22.15%

Percent of the sources are human = 161 / 623 = 25.84%

Percent of the sources are terrestrial mammal = 56 / 623 = 8.99%

b.

The percentage of locations to be calculated by finding the ratio between the location totals and the species total.

Percent of the location Chorro Creek = 241 / 623 = 38.68%

Percent of the location Los Osos Creek = 256 / 623 = 41.09%

Percent of the location Baywood Seeps = 126 / 623 = 20.22%

c.

The rate of contamination from farm animals for the three locations can be calculated by finding the ratio between the fecel contamination by farm animals at respective locations and the respective location totals.

Rate of contamination by farm animals at Chorro Creek = 106 / 241 = 43.98%

Rate of contamination by farm animals at Los Osos Creek = 32 / 256 = 12.5%

Rate of contamination by farm animals at Baywood Seeps = 0 / 126 = 0%

Among the above three rate, the highest rate is 43.98%. So the Chorro Creek contains the highest rate of contamination from farm animals.

d.

The rate of contamination from humans for the three locations can be calculated by finding the ratio between the fecel contamination by human at respective locations and the respective location totals.

Rate of contamination by humans at Chorro Creek = 38 / 241 = 15.77%

Rate of contamination by humans at Los Osos Creek = 63 / 256 = 24.61%

Rate of contamination by humans at Baywood Seeps = 60 / 126 = 47.62%

Among the above three rate, the highest rate is 47.62%. So the Baywood Seeps contains the highest rate of contamination from humans.

e.

The rates of contamination by source in Los Osos Creek can be determined by finding the ratio between the respective sources at Los Osos Creek and the species total at Los Osos Creek.

Rate of contamination by birds at Los Osos Creek = 79 / 256 = 30.86%

Rate of contamination by Domestic pets at Los Osos Creek = 56 / 256 = 21.88%

Rate of contamination by farm animals at Los Osos Creek = 32 / 256 = 12.5%

Rate of contamination by humans at Los Osos Creek = 63 / 256 = 24.61%

Rate of contamination by Terrestrial mammals at Los Osos Creek = 26 / 256 = 10.16%

Put in decreasing order.

30.86% . 24.61%, 21.88%, 12.5%, 10.16%

Birds, Humans, Domestic pets, Farm animals, Terrestrial mammals.

f.

The percent of samples with terrestrial mammal E.coli contamination in Chorro Creek can be determined by finding the ratio between the terrestrial mammals at respective locations and the total of terrestrial mammals.

Percent of samples with terrestrial mammal E.coli contamination in Chorro Creek = 22 / 56 = 39.39%

Percent of samples with terrestrial mammal E.coli contamination in Los Osos Creek = 26 / 56 = 46.43%

Percent of samples with terrestrial mammal E.coli contamination in Baywood Seeps = 8 / 56 = 14.29%

The percent of samples with terrestrial mammal E.coli contamination in Chorro Creek is 39.39%